OpenStudy (anonymous):

Intensity, I, of light received at a source varies inversely as the square of the distance, d, from the source. If the light intensity is 91 ft-candles at 18ft, find the intensity at 9ft. I=k/d^2

OpenStudy (amistre64):

91 = k/(18^2) 91(18^2) = k l = 91(18^2)/(9^2)

OpenStudy (amistre64):

l = 91(18/9)^2 l = 91(2)^2 l = 91(4)